OST Or smpping 22. Firebrick Companys budgeted sales were 10,000 units at $200 per unit. Actual sales were 2,250 units at $210 per unit Firebricks sales price variance was: A) $ 34,000 (U) B) $ 22,500 (F) C) $ 90,000 (F) D) S 45,000 (F)
Answer #1

Answer is $ 22,500(F)

Sales Price Variance (SPV) = Actual Total Price of Goods Sold (-) Standard Total Price of Actual Sales

As,.Goods Sold or Actual Sales in units is 2250 Units.

SPV = (Actual Sales Price * No. of Units Sold) - (Standard Price * Total Sales)

SPV = 210*2250 (-) 200*2250

SPV = 2250(210 - 200)

SPV = 2250*10

As Actual Sales Price is more than Standard Sales Price for a given quantity sold thus its is Favourable Variance.